OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -2x - y < -6 Choose one answer. a. Dashed line with shading on the side that includes the origin. b. Solid line with shading on the side that does not include the origin. c. Dashed line with shading on the side that does not include the origin. d. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

so it is less than which means dashed?

OpenStudy (danjs):

put in the form y > .... first

OpenStudy (anonymous):

oh alright

OpenStudy (danjs):

y > -2x + 6

OpenStudy (anonymous):

dashed still?

OpenStudy (danjs):

yes, \[\gt~or ~\lt~~dashed\] \[\ge~or~\le~~solid line\]

OpenStudy (danjs):

just remember it can equal the line, so solid line means include those points on the line

OpenStudy (anonymous):

oh alright

OpenStudy (danjs):

now you can just pick 2 points and connect them, x=0 and x=1

OpenStudy (danjs):

When x=0, y = 6, point (x,y) = (0,6) When X=1, y = -2 + 6 = 4 . point (x,y) = (1,4)

OpenStudy (anonymous):

this does not include the origin correct?

OpenStudy (danjs):



OpenStudy (anonymous):

http://gyazo.com/9ad5ae8bf49066d458c9301d46619cf4

OpenStudy (danjs):

Since it is "Less than" you shade underneath the line...

OpenStudy (danjs):

Greater than, you would shade above

OpenStudy (anonymous):

so above

OpenStudy (anonymous):

you said earlier the equation was different. i showed it in my link

OpenStudy (danjs):

yeah i just noticed i mistyped,... right y > -2x + 6 , greater than , shade above the line

OpenStudy (anonymous):

are you sure its dashed line for greater than or less than?

OpenStudy (anonymous):

i would think it would be solid in that case

OpenStudy (danjs):

yes greater than or less than, dashed line, it doesnt include the line,

OpenStudy (anonymous):

so the answer to this one was c

OpenStudy (danjs):

yes, another quick thing you could do for these types probs, is simply test the origin

OpenStudy (danjs):

x = 0, y = 0 -2*0 - 0 < - 6 0 < 6 False, the origin is not part of the solution/shaded area

OpenStudy (danjs):

and it is dotted because of > or <

OpenStudy (anonymous):

could i check another equation?

OpenStudy (danjs):

sure, we can do it the quick way

OpenStudy (anonymous):

just wanna make sure im doing it right i got the answer already i think

OpenStudy (anonymous):

http://gyazo.com/f91fa996a6ecec7fd818376c3715b8ec

OpenStudy (anonymous):

a. 2x - 5y -10 b. x - y > 3 c. 2x - y < -4 d. x + 2y -8

OpenStudy (anonymous):

i chose d

OpenStudy (danjs):

what is the sign in d

OpenStudy (anonymous):

greater than or equal to

OpenStudy (danjs):

ok, that part is right, it is a solid line and shaded above > or = to

OpenStudy (danjs):

now test a couple points on the line... When x =0, y=-4 on the graph, test in answer d

OpenStudy (danjs):

0 + 2y = -8 y = -4 , so far so good, now test one more point

OpenStudy (danjs):

When y=0, x = -8 on the graph.. x + 2(0) = -8 True also, choice d works

OpenStudy (anonymous):

Alright thanks for the help

OpenStudy (danjs):

no prob, good luck